Two resistors, 58.0 Ohms and 58.0 Ohms , are connected in parallel. The current through the second resistor is 4.40 A.

(a) What is the current in the first resistor?_______A
(b) What is the total power supplied to the two resistors?________W
1
Expert's answer
2012-02-14T09:44:25-0500
Two resistors, 58.0 Ohms& and 58.0 Ohms , are connected in parallel. The current through the second resistor is 4.40 A.
&
(a) What is the current in the first resistor?________A

Answer: 4.40 A

If two resistors with the same resistances are connected in parallel, the current through the first resistor is the same as the current through the second resistor.

(b) What is the total power supplied to the two resistors?________W

Answer: 2245.76 W

P1 = P2 = RI² = 58*4.4² = 1122.88 W
P = P1+P2 = 2*1122.88 = 2245.76 W